OpenStudy (anonymous):

The polynomial 3x^2+bx-4 has x-a as a factor. where a is not equal to 0 express b in terms of a

ganeshie8 (ganeshie8):

hint : if (x-a) is a factor of f(x), then f(a) = 0

OpenStudy (anonymous):

@ganeshie8 thanks! but i srsly dont know how to answer this tho. i need help T.T

ganeshie8 (ganeshie8):

since x-a is a factor of given polynomial(say f(x)), f(a) must equal 0

ganeshie8 (ganeshie8):

plugin x = a and set it equal to 0

ganeshie8 (ganeshie8):

3x^2+bx-4 plugin x = a, what do u get ?

OpenStudy (anonymous):

3a^2 + ba - 4 = 0. am i right?

ganeshie8 (ganeshie8):

Excellent ! next just isolate `b`

OpenStudy (anonymous):

so... ba = -3a^2 + 4 and then, b = -3a^2 + 4/a. is the answer b = -3a^2 + 4?

ganeshie8 (ganeshie8):

yes 3a^2 + ba - 4 = 0 ba = -3a^2 + 4 b = (-3a^2+4)/a

OpenStudy (anonymous):

-3a^2 + 4/a is the answer not -3a^2 + 4?

OpenStudy (anonymous):

cant i cancel a in the denominator by the a in 3a^2?

ganeshie8 (ganeshie8):

thats a good observation ! yes you can

ganeshie8 (ganeshie8):

b = (-3a^2+4)/a = -3a^2/a + 4/a = -3a + 4/a

OpenStudy (anonymous):

thank you so much!!!

ganeshie8 (ganeshie8):

np :)